OpenStudy (anonymous):

Convert each measure to degrees. Round answers to the nearest hundredth if needed. a.) pi/9 b.) -2pi/3

Nnesha (nnesha):

formula is \[\huge\rm r \times \frac{ 180 }{ \pi }\] r= radian

OpenStudy (anonymous):

2π radians = 360° 1 radian = 360°/2π = 180°/π 1.6 radian = 1.6*180/π use your calculator or a pencil and paper to calculate 1.6*180/π and you'll have your answer.

OpenStudy (anonymous):

wait ok you just lost me? @ElmosGuts my answer for which one?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ok i got 91.673? @ElmosGuts

Nnesha (nnesha):

\[\huge\rm \frac{ \pi }{ 9 } \times \frac{ 180 }{ \pi}= ?\] to convert radian to degree you should apply this formula radian multiply by 180/pi

OpenStudy (anonymous):

so 20? @Nnesha

Nnesha (nnesha):

yes

OpenStudy (anonymous):

so the second one i got -120 is that right?

Nnesha (nnesha):

yeah :-)

OpenStudy (anonymous):

thank you so much!(:

Nnesha (nnesha):

np